Ramon rented a sprayer and a generator. On his first job, he used the sprayer for 6 hours and the generator for 6 hours at a total cost of $90. On his second job, he used the sprayer for 4 hours and the generator for 8 hours at a total cost of $100. What was the hourly cost of each piece of equipment? How do I solve?
Answer by Boreal(15235)

S=sprayer
G=generator
6S+6G=90
4S+8G=100
multiply top by 2 and bottom by -3 to eliminate S
12S+12G=180
-12S-24G=-300
-12G=-120
G=$10
Therefore, S=$5 per hour